# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v2

@@ 4.5 Aerodynamic drag and lift coefficients
Accurate and fast predictions of global quantities, such as drag and lift coefficients, are crucial for designing efficient aerodynamic geometries. We evaluate the quality of the surface-level predictions of our model by computing these integrated aerodynamic quantities. Figure 6 shows prediction and reference values for drag and lift coefficients on DrivAerML. AB-UPT obtains accurate predictions as indicated by the high coefficient of determination ($R^{2}$) score. Additionally, such integrated quantities require high-resolution predictions (see Figure 1), which AB-UPT can effortlessly produce via anchor attention, despite being trained on only a small subset (e. g., 16 thousand) surface points.
@@ 4.6 Training from CAD
Table 5: Neural simulation without using the CFD simulation mesh on AhmedML. AB-UPT can produce accurate results even in the absence of the CFD simulation mesh, which is typically costly to obtain. Numerical simulations (e. g., HRLES) would not converge or produce unrealistic results without the CFD simulation mesh. Note that HRLES with the CFD simulation mesh is used as ground truth (i.e., 0.00 error).
Relative L2 error ($\downarrow$) Meshing
$\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ Surface mesh Volume mesh
HRLES 0.00 0.00 0.00 CFD simulation mesh CFD simulation mesh
HRLES N.A. N.A. N.A. Geometry mesh (CAD) Regular grid
AB-UPT 3.14 2.07 6.34 CFD simulation mesh CFD simulation mesh
AB-UPT 3.62 2.56 7.43 Geometry mesh (CAD) Regular grid
@@ 4.7 Divergence-free vorticity
As discussed in Sections 3.2 and 3.3, the conditional field property of AB-UPT allows for defining vorticity field predictions that are divergence-free by design. We train the model in two stages: in the pretraining phase, we learn a model that predicts all surface and volume fields except the vorticity, using the same model configuration and hyperparameters as in the previous experiments; in the finetuning phase, we learn all fields including the derived vorticity field (cf. Section 3.3). Furthermore, we use float32 precision for finetuning, and we train with a smaller initial learning rate and fewer training steps: starting from $2.0\cdot 10^{-5}$, the learning rate is decayed to $1.0\cdot 10^{-6}$ via the cosine annealing schedule. As discussed in Section 3.3, after correcting for the scaling of the velocity field due to normalization, the predicted vorticity is already in physics space. To bring the predictions to a similar scale as other predicted fields, we transform both the predictions and targets as follows: first divide by the field’s standard deviation; then apply a sign-preserving square-root to the vector’s magnitude. This is equivalent to scaling the individual loss terms. All other fields are simply normalized as in all other experiments, subtracting the mean and dividing by the standard deviation (c. f. Section B.4.1).
@@
Table 6: Evaluation of AB-UPT with the divergence-free vorticity formulation (c. f. Section 3.3). For each dataset, the first 3 columns per datasets show the relative L2/L1 test errors (in %) of surface pressure $\mathbf{p}_{s}$, volume velocity $\mathbf{u}$ and volume vorticity $\bm{\omega}$. The last column shows the mean absolute value of the divergence of the predicted vorticity field $\nabla\cdot\bm{\omega}$ (in $m^{-1}s^{-1}$), where the average is taken over simulations and samples of the field.
@@
$\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ $\nabla\cdot\bm{\omega}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ $\nabla\cdot\bm{\omega}$
@@ 5 Limitations and future work
AB-UPT employs expressive self-attention among anchor tokens, which exhibits quadratic complexity. We show that on high-fidelity CFD simulations a small set of anchor tokens is sufficient to obtain good results on currently available public datasets. However using more anchor tokens can improve accuracy further, leading to (quadratically) increasing computational demands. Contrary, the computational demands of models with linear complexity would only increase linearly.
However, we did not manage to scale linear attention models to obtain higher accuracies than AB-UPT, even when using twice the number of inputs (see Figure 5). Additionally, the trend of the loss curves in Figure 5 suggests that the benefit of scaling the number of inputs saturates. As there is still a sizable gap between the AB-UPT test loss and linear attention test loss (i.e., Transolver), it is unlikely that linear attention models would obtain better accuracies than AB-UPT, even when trained on a large number of inputs. Nevertheless, for an extremely large number of anchor tokens, compute demands would explode, and AB-UPT would need some kind of linear attention mechanisms.
@@
AB-UPT is a general-purpose architecture, which we expect to be applicable to many use cases beyond automotive CFD. We chose automotive CFD as a problem setting as there exist public high-fidelity datasets, which serve as a good starting point where models such as UPT and Transolver start to break as they can not model the complex dynamics of these datasets. However, as AB-UPT is a natural successor of UPT, we expect it to perform well on many simulation types, as is the case for UPT (e.g., transient fluid simulations (Alkin et al., 2024a) or particle and particle-fluid multi-physics simulations (Alkin et al., 2024b)).
6 Conclusion and outlook
@@ Appendix A Extended results
A.1 Full results: benchmarking AB-UPT against other neural surrogate models
Table 7: Relative L2 errors (in %) of surface pressure $\mathbf{p}_{s}$, volume velocity $\mathbf{u}$, volume vorticity $\bm{\omega}$, wall shear stress $\bm{\tau}$, and volume pressure $\mathbf{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for ShapeNetCar, AhmedML, and DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. For ShapeNetCar, only surface pressure and volume velocity are available.
ShapeNetCar AhmedML DrivAerML
$\mathbf{p}_{s}$ $\mathbf{u}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ $\bm{\tau}$ $\mathbf{p}_{v}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ $\bm{\tau}$ $\mathbf{p}_{v}$
GINO 13.28 2.53 7.90 6.23 71.81 8.18 8.10 3.03 40.58 98.5 21.71 44.90
LNO 9.05 2.29 12.91 7.56 72.05 11.44 8.47 20.51 23.27 173.0 36.44 27.02
UPT 5.60 1.40 3.90 2.62 13.69 5.52 3.01 7.85 9.02 189.6 12.70 10.54
OFormer 7.05 1.61 3.55 2.21 7.78 4.10 2.27 4.48 6.64 60.4 8.92 7.10
Transolver 6.50 1.53 3.66 2.19 8.12 4.00 2.16 4.81 6.78 38.4 8.95 7.74
Transformer 5.01 1.24 3.41 2.09 6.76 4.03 2.16 4.35 6.21 47.9 8.26 6.27
AB-UPT 4.81 1.18 3.00 1.88 6.45 3.88 1.98 3.82 5.93 35.1 7.29 6.08
A.2 Benchmarking AB-UPT against DoMINO
Table 8: Relative L2 errors (in %) of surface pressure $\mathbf{p}_{s}$, wall shear stress $\bm{\tau}$ per $x,y,z$, dimension, volume velocity $\mathbf{u}$ per $x,y,z$ dimension, and volume pressure $\mathbf{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. The results for DoMINO are taken from (Ranade et al., 2025)*.
@@ A.2 Benchmarking AB-UPT against DoMINO
A.3 Analysis: Anchor Attention

### v3

@@ 4.5 Aerodynamic drag and lift coefficients
Accurate and fast predictions of global quantities, such as drag and lift coefficients, are crucial for designing efficient aerodynamic geometries. We evaluate the quality of the surface-level predictions of our model by computing these integrated aerodynamic quantities by predicting pressure $\mathbf{p}$ and wallshearstress $\mathbf{\bm{\tau}}$ for each of the 9 million surface mesh cells using 16 thousand anchor tokens (employing chunked inference). These predictions are used as $\bm{p}$ and $\mathbf{\bm{\tau}}$ in Equation 2 to calculate the force acting on the car $\bm{F}$. We then plug in $\bm{F}$ into Equations 3 to obtain the drag coefficient $C_{d}$ and the lift coefficient $C_{l}$. Figure 6 shows prediction and reference values for drag and lift coefficients on DrivAerML. AB-UPT obtains accurate predictions as indicated by the high coefficient of determination ($R^{2}$) score. Additionally, such integrated quantities require high-resolution predictions (see Figure 1), which AB-UPT can effortlessly produce via anchor attention, despite being trained on only a small subset (e. g., 16 thousand) surface points.
@@ 4.6 Training from CAD
Table 5: Relative L2 errors (in %) of neural simulation without using the CFD simulation mesh. AB-UPT can produce accurate results even in the absence of the CFD simulation mesh, which is typically costly to obtain. Numerical simulations (e. g., HRLES) would not converge or produce unrealistic results without the CFD simulation mesh. Note that HRLES with the CFD simulation mesh is used as ground truth (i.e., 0.00 error).
AhmedML DrivAerML Meshing
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ Surface mesh Volume mesh
HRLES 0.00 0.00 0.00 0.00 0.00 0.0 CFD simulation mesh CFD simulation mesh
HRLES N.A. N.A. N.A. N.A. N.A. N.A. Geometry mesh (CAD) Regular grid
AB-UPT 3.01 1.90 6.52 3.82 5.93 35.1 CFD simulation mesh CFD simulation mesh
AB-UPT 3.70 2.50 7.59 4.14 6.64 76.5 Geometry mesh (CAD) Regular grid
@@ 4.7 Divergence-free vorticity
As discussed in Sections 3.2 and 3.3, the conditional field property of AB-UPT allows for defining vorticity field predictions that are divergence-free by design. We train the model in two stages: in the pretraining phase, we learn a model that predicts all surface and volume fields except the vorticity, using the same model configuration and hyperparameters as in the previous experiments; in the finetuning phase, we learn all fields including the derived vorticity field (cf. Section 3.3). Furthermore, we use float32 precision for finetuning, and we train with a smaller initial learning rate and fewer training steps: starting from $2e^{-5}$, the learning rate is decayed to $1e^{-6}$ via the cosine annealing schedule. As discussed in Section 3.3, after correcting for the scaling of the velocity field due to normalization, the predicted vorticity is already in physics space. To bring the predictions to a similar scale as other predicted fields, we transform both the predictions and targets as follows: first divide by the field’s standard deviation; then apply a sign-preserving square-root to the vector’s magnitude. This is equivalent to scaling the individual loss terms. All other fields are simply normalized as in all other experiments, subtracting the mean and dividing by the standard deviation (c. f. Section B.4.1).
@@
Table 6: Evaluation of AB-UPT with the divergence-free vorticity formulation (c. f. Section 3.3). For each dataset, the first 3 columns per datasets show the relative L2/L1 test errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$ and volume vorticity $\bm{\omega}$. The last column shows the mean absolute value of the divergence of the predicted vorticity field $\nabla\cdot\bm{\omega}$ (in $m^{-1}s^{-1}$), where the average is taken over simulations and samples of the field.
@@
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\nabla\cdot\bm{\omega}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\nabla\cdot\bm{\omega}$
@@
4.8 Scaling AB-UPT
We investigate three axes of scaling (as visualized in Figure 2(a)): problem-scale, model-scale, and data-scale. In every study, one scale is varied, whereas the other two scales are fixed. By default, we use 16K/16K surface/volume anchor tokens (problem-scale), 12 blocks with a dimension of 192 (model-scale), and all 400 training simulations of DrivAerML (data-scale), which is the setting used throughout the previous sections.
Figure 7: Left and middle: Scaling the number of anchor tokens during inference only. Models are trained with varying anchor tokens (as indicated by their starting point and label). These models are then evaluated with an increasing number of anchor tokens but without any retraining (as indicated by the x-axis). Different colors correspond to different numbers of anchor tokens used during training. Increased number of anchor tokens used in training yields performance improvements; increased number of anchor tokens during inference yields similar accuracies. Right: Summary of the left and middle plot where the number of inference anchors is fixed at 262K (i.e., the best obtained loss) and the x-axis varies the number of training anchors instead. The performance saturation hints at the problem scale of the DrivAerML dataset, i.e., certain amounts of anchor tokens are enough to capture the intricacies of the dataset.
Problem-scale
We train AB-UPT models with various numbers of anchor tokens (4K to 131K) and additionally increase the number of anchor tokens further during inference only (up until 262K), without any retraining. By increasing the number of anchor tokens, we increase coverage of the 150M simulation mesh cells and observe saturating effects when increasing the number of anchor tokens beyond a certain point. Figure 7 shows results thereof. The left and middle plots show that AB-UPT benefits from an increased number of anchor tokens in inference time, but the best models are obtained by also training with more anchor tokens. As the number of anchor tokens increases, performance gains saturate (e.g., doubling training anchors from 4K to 8K decreases the test loss by 8%, doubling anchors from 32K to 64K decreases test loss by only 2%).
Interestingly enough, a model trained with 131K anchor tokens already shows signs of overfitting, where the train loss still decreases (is better than with 65K tokens) but test loss increases (is worse than with 65K tokens), as shown on the right side of Figure 7. This hints at the underlying problem-scale of DrivAerML, i.e., 65K tokens in training are sufficient to capture the underlying dynamics of the problem. Notably, this is three orders of magnitude less than the 150M mesh cells used in the classical CFD simulation.
Model-scale
We investigate two ways to scale model size: increasing the dimension of the model and increasing the depth of the model and present results in Figure 8. As expected, increasing model size improves the train loss. The test loss also increases until overfitting becomes a problem.
Figure 8: Left and middle: Model scaling by scaling either the dimension (using a depth of 12) or the depth (using a dimension of 192). Right: Scaling dataset size.
Data-scale
As obtaining more high-resolution CFD data is costly (roughly \$2000 per simulation), we investigate scaling the dataset size by training on subsets of the full dataset. Figure 8 shows a clear trend where more data leads to better generalization performance.
@@ 5 Limitations and future work
AB-UPT employs expressive self-attention among anchor tokens, which exhibits quadratic complexity. We show that on high-fidelity CFD simulations a small set of anchor tokens is sufficient to obtain good results on currently available public datasets. However, using more anchor tokens can improve accuracy further, leading to (quadratically) increasing computational demands. Contrary, the computational demands of models with linear complexity would only increase linearly.
However, we did not manage to scale linear attention models to obtain higher accuracies than AB-UPT, even when using an order of magnitude more inputs (see Figure 5). Additionally, the trend of the loss curves in Figure 5 suggests that the benefit of scaling the number of inputs saturates. As there is still a sizable gap between the AB-UPT test loss and linear attention test loss (i.e., Transolver), it is unlikely that linear attention models would obtain better accuracies than AB-UPT, even when trained on a large number of inputs. Nevertheless, for an extremely large number of anchor tokens, compute demands would explode, and AB-UPT would need some kind of linear attention mechanisms. For example, ball attention (Zhdanov et al., 2025) is a promising attention mechanism with linear complexity that could be integrated into AB-UPT.
@@
AB-UPT is a general-purpose architecture, which we expect to be applicable to many use cases beyond automotive CFD. We chose automotive CFD as a problem setting due to existing publicly available high-fidelity datasets. However, as AB-UPT is a natural successor of UPT, we expect it to perform well on many simulation types, as is the case for UPT (e.g., transient fluid simulations (Alkin et al., 2024a) or particle and particle-fluid multi-physics simulations (Alkin et al., 2024b)).
High-fidelity CFD simulations
The ever-lasting paradigm that a surrogate model can only achieve the accuracy of its underlying simulation is particularly valid in CFD. The chosen turbulence modeling approach (e.g., RANS, LES, HRLES) and the adopted meshing strategy critically determine the fidelity and accuracy of the simulation results. This is especially pronounced in external automotive aerodynamics, where capturing complex near-wall flow behavior holds significant challenges and remains a key area for methodological advancement. In this context, CFD approaches such as the Lattice Boltzmann Method (Chen & Doolen, 1998) are of growing interest due to their inherently parallelizable structure, which facilitates efficient deployment on GPU architectures. Accelerated simulation techniques enable the resolution of finer-scale flow structures and near-wall phenomena, thereby improving the physical realism of the computed flow fields. Enhancing the underlying simulation quality directly benefits the machine learning-based surrogate models, leading to higher accuracy, better generalization, and increased reliability.
6 Conclusion
@@ Appendix A Extended results
A.1 Training without query tokens
As anchor attention shares its weights for the anchors and queries, one would ideally train a model without employing a loss on the queries, as propagating query tokens during training imposes a runtime overhead. However, there is a conceptual difference between how anchors interact with each other and how queries interact with anchors. Namely, anchor tokens attend to themselves and can influence the representation of other anchor tokens. While this is conceptually very similar to how query tokens are propagated, it imposes a train-test discrepancy if no query tokens are used during training. We investigate the influence of this discrepancy by training models where the loss is calculated on (i) only anchor tokens (ii) only query tokens (iii) anchor and query tokens. We then evaluate these models using a variable number of query tokens to investigate the generalization capabilities of anchor attention when trained with a loss on the anchor tokens. Table 7 shows that the train-test discrepancy from (i) leads to a slight decrease in performance, which is not present in (ii) or (iii). We use (i) throughout the main paper due to its computational efficiency.
Table 7: The train-test discrepancy of training with only anchor tokens (i), leads to a slight increase in loss when query tokens are used in inference. Including queries during training (ii) and (iii) leads to almost the same loss on anchors and queries. Training with a loss on anchors and queries (iii) obtains the best performance. Relative increase denotes the increase from the anchor loss to the query loss. We use DrivAerML with 16K anchors/queries for this study. Train speed denotes the runtime for 100 updates with batchsize 1.
Training with loss on Loss when evaluated on
Anchors Queries Anchors ($\downarrow$) Queries ($\downarrow$) Relative increase ($\downarrow$) Train speed
(i) 0.13574 0.13624 $+0.37\%$ 11.5s
(ii) 0.13634 0.13632 $\mathbf{-0.01\%}$ 19.4s
(iii) 0.13462 0.13468 $\mathbf{+0.04\%}$ 19.4s
Adding increasingly many query tokens adds computational costs and poses a trade-off. Figure 9 visualizes this trade-off, where an increasing number of queries improves the performance of the predictions from the query tokens. When employing a combination of anchor and query losses (iii), fewer query tokens can be used.
Figure 9: Performance of predictions from query tokens on DrivAerML when trained with various number of query tokens. The number of anchor tokens is fixed at 16K. The values at 16K query tokens correspond to (ii) and (iii) of Table 7 and the value at 0K corresponds to (i). Train time is measured as runtime for 100 updates with batchsize 1.
A.2 Model design on DrivAerML
Table 8: Impact of model design on DrivAerML. The design steps follow a similar pattern as in Table 3 where the impact of step-by-step model design choices are shown on AhmedML.
L2 error ($\downarrow$) Neural field Mesh independent Train speed
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
baseline UPT 12.67 16.65 10424 7.0
macro + large decoder 7.44 8.74 90.2 14.7
design + decoder-only with self-attn 4.35 6.21 47.9 14.9
re-add + anchor attention 4.35 6.21 47.9 14.9
properties + decouple geometry encoding 4.12 5.90 33.7 16.1
micro + split surface/volume branch 4.23 6.09 32.6 11.0
design + cross-branch interactions 4.24 5.97 35.8 11.0
+ branch-specific decoders 3.82 5.93 35.1 11.0
A.3 Benchmarking AB-UPT against DoMINO
Table 9: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, wall shear stress $\bm{\tau}$ per $x,y,z$, dimension, volume velocity $\bm{u}$ per $x,y,z$ dimension, and volume pressure $\bm{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. The results for DoMINO are taken from (Ranade et al., 2025)*.
@@ A.3 Benchmarking AB-UPT against DoMINO
A.4 Full results: benchmarking AB-UPT against other neural surrogate models
Table 10: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, volume vorticity $\bm{\omega}$, wall shear stress $\bm{\tau}$, and volume pressure $\bm{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for ShapeNet-Car, AhmedML, and DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. For ShapeNet-Car, only surface pressure and volume velocity are available.
ShapeNet-Car AhmedML DrivAerML
$\bm{p}_{s}$ $\bm{u}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{\tau}$ $\bm{p}_{v}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{\tau}$ $\bm{p}_{v}$
GINO 13.28 2.53 7.90 6.23 71.81 8.18 8.10 3.03 40.58 98.5 21.71 44.90
LNO 9.05 2.29 12.91 7.56 72.05 11.44 8.47 20.51 23.27 173.0 36.44 27.02
UPT 5.60 1.40 4.25 2.73 15.03 5.80 3.10 7.44 8.74 90.2 12.93 10.05
OFormer 7.05 1.61 3.55 2.21 7.78 4.10 2.27 4.48 6.64 60.4 8.92 7.10
Transolver 6.46 1.62 3.45 2.05 8.22 4.00 2.16 4.81 6.78 38.4 8.95 7.74
Transformer 4.86 1.17 3.41 2.09 6.76 4.03 2.16 4.35 6.21 47.9 8.26 6.27
AB-UPT 4.81 1.16 3.01 1.90 6.52 3.88 1.98 3.82 5.93 35.1 7.29 6.08
A.5 Benchmarking AB-UPT against Erwin
Erwin (Zhdanov et al., 2025) is a promising linear complexity transformer variant that employs self-attention within local regions, so-called balls, with efficient memory structures to enable computational efficiency on arbitrary pointclouds. This changes the runtime complexity of the attention from $\mathcal{O}(N^{2})$ to $\mathcal{O}((N/B)*B^{2})$ where $N$ is the number of inputs, $N/B$ is the number of balls and $B$ is the ball size, i.e., the number of tokens within a single ball. If $B=N$, it reduces to full self-attention.
As similar local attention variants designed for regular grids (Liu et al., 2021) have been very successful employed in neural simulation (e.g., weather modeling (Bodnar et al., 2025)), we consider ball attention a promising direction with one of the best runtime-accuracy tradeoff among linear complexity transformers. Combinations of AB-UPT and ball attention are interesting future directions, where anchor attention could be extended to support local ball structures. Similarly the cross-attention components of AB-UPT could be adapted to cross ball attention.
However, as shown in the main paper (e.g., Figure 5), training full quadratic self-attention models is perfectly feasible even on the problem-scale of DrivAerML employing >100M mesh cells for its HRLES simulations. Consequently, it is unlikely that the local attention obtains better accuracy than full self-attention but if runtime is a requirement, local attention could speedup the model while largely preserving accuracy. In the following, we aim to provide insights into this trade-off by adapting ball attention to our joint surface/volume variable modeling setting. This adaption, termed Erwin (reimpl.), uses an (i) isotropic architecture, i.e., there are no up/downsampling operations in the model, (ii) does not use local attention biases and (iii) uses two single layer message passing modules for surface/volume encoding respectively, instead of three layers of message passing as in the original architecture. (i) Is used to have a roughly amount of parameter count to AB-UPT. (ii) Speeds up computations and keeps the vanilla self-attention computation instead of biasing it. (iii) Aligns message passing complexity with that of supernode pooling where multi-layer message passing would become extremely costly with large number of inputs. After encoding surface and volume points, we concatenate them and process them with 12 ball attention blocks with a dimension of 192 (same setting as for the other transformer-based models). We do not add rotational embeddings (RoPE) (Su et al., 2024) to ball attention blocks as the original implementation [^3] https://github.com/maxxxzdn/erwin does not support it, which could yield further performance improvements. Also hierarchical designs could yield further improvements, which we do not consider as there is a large design space for hierarchical designs and they are hard to directly comparable in terms of FLOPS/parameter counts.
We first compare on ShapeNet-Car where we can easily study variable ball size counts. Table 11 shows that Erwin is a very promising linear complexity transformer variant performing best on ShapeNet-Car among the considered ones in this paper. However, quadratic complexity transformers obtain better accuracies.
Table 11: Benchmarking our adaption of Erwin to jointly model surface and volume data on ShapeNet-Car. Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, and volume velocity $\bm{u}$. Erwin (reimpl.) with ball size 8192 would be roughly equivalent to the Transformer row.
Ball size $\bm{p}_{s}$ $\bm{u}$
Linear complexity transformers
LNO - 9.05 2.29
OFormer - 7.05 1.60
Transolver - 6.46 1.62
Erwin (reimpl.) 512 5.46 1.37
Erwin (reimpl.) 1024 5.38 1.37
Erwin (reimpl.) 2048 5.35 1.27
Erwin (reimpl.) 4096 5.31 1.25
Quadratic complexity transformers
Transformer 8192 4.86 1.17
AB-UPT 4096 4.81 1.16
Next, we compare our reimplementation of Erwin on DrivAerML in the default setting considered in the main paper (16K/16K surface/volume points) and show results in Figure 10. We observe that ball attention imposes an overhead, most likely from memory restructuring operations to structure tokens for efficient attention computations. This overhead becomes negligible with more inputs and ball attention becomes faster than (branched) self-attention. As there is a significant accuracy gap between AB-UPT trained on 16K/16K points and Erwin (reimpl.) trained on 16K/16K points we hypothesize that more sophisticated training protocols are necessary to train ball tree attention models to sufficient quality (e.g., hierarchical architectures, different ball size to input token ratios, …). As the design space thereof is large and fair comparisons to AB-UPT become more and more difficult, we leave exploration of this direction to future work.
Figure 10: Left: Test loss of Erwin (reimpl.) trained on DrivAerML. Right: Runtime comparison of plain self-attention and ball attention blocks. Train time is measured in seconds per 100 updates with batchsize 1 with 12 blocks of dimension 192.
Finally, we compare AB-UPT against the original Erwin architecture (including three layer message passing and distance-based attention biases) in the setting used in the original paper (Zhdanov et al., 2025). Table 12 confirms that our reimplementation obtains sensible results and that AB-UPT also performs best with this protocol.
Table 12: Comparing AB-UPT and Erwin (reimpl.) against Erwin on ShapeNet-Car surface pressure only prediction (no volume data is used). We use a similar protocol to Erwin (300 epochs with early stopping, 700/189 train/test split from (Alkin et al., 2024a), median of 5 runs). Results for Erwin are taken from (Zhdanov et al., 2025). Note that Erwin (reimpl.) uses 12 blocks of dimension 192 and consequently has more parameters than Erwin-M. Ball size is 256.
$\bm{p}_{s}$ MSE
Erwin-S 15.85
Erwin-M 15.43
Erwin (reimpl.) 15.17
AB-UPT 13.25
A.6 Anchor attention versus k-NN interpolation